# Generalized Schalkwijk-Kailath Coding for Autoregressive Gaussian Channels

### v1

@@ 4 The SK(2) Coding Scheme
which implies that $I_{n}+UU^{\ast}$ is positive definite and thus invertible. Then, we use the Woodbury matrix identity [6] to establish
$(I_{n}+UU^{\ast})^{-1}$ $=I_{n}-U(I_{m}+U^{\ast}U)^{-1}U^{\ast}$
$=I_{n}-UA^{-1}U^{\ast}$
@@
$\widetilde{Y}_{n}=d_{n}U+W_{n}$
for $n\geq p+1$, where
@@
Thus, the mutual information between $V_{1}^{n}$ and $Y_{1}^{n}$ for $n\geq p+1$ is given by
@@
$\geq\frac{1}{n}I(U;\widetilde{Y}_{p+1}^{n})$
$=-\frac{1}{2n}\log\mathbb{E}\left[\left|U-\mathbb{E}[U|\widetilde{Y}_{p+1}^{n}\right|^{2}\right].$
Let
$D_{n}=\begin{bmatrix}d_{1}&d_{2}&...&d_{n}\end{bmatrix}^{T}.$
From (53) we obtain
$\mathbb{E}[|U-\mathbb{E}[U|\widetilde{Y}_{1}^{n}|^{2}]$ $=\frac{1}{1+\mathrm{Tr}(D_{n}D_{n}^{T})}$ (55)
$=\frac{1}{1+D_{n}^{T}D_{n}}$
$=\bigl[1+\frac{L^{2}(\gamma^{-1})}{\gamma^{2}-1}(\gamma^{2n}-1)\bigr]^{-1}.$
Substituting this into (54) and letting $n\to\infty$ yields (12), as desired.
Moreover, following a similar approach as in (26), we observe that for $n\geq p+2$
@@
$\leq\mathbb{E}[|V_{n}-\mathbb{E}[V_{n}|\widetilde{Y}_{p+1}^{n-1}]|^{2}]$
$=\gamma^{2(n-1)}\mathbb{E}[|U-\mathbb{E}[U|\widetilde{Y}_{p+1}^{n-1}]|^{2}],$
where the inequality follows because conditioning on a larger $\sigma$-algebra reduces the estimation error. Combining (55) with (56) yields
[…]
@@
Finally, it follows from (3) that the power constraint (13) is satisfied, which completes the proof.

### v2

@@ 4 The SK(2) Coding Scheme
which shows that $S+UU^{\ast}$ is positive-definite and hence invertible. Then, we use the Woodbury matrix identity [6] to establish
$(S+UU^{\ast})^{-1}$ $=S^{-1}-S^{-1}U(I_{m}+U^{\ast}S^{-1}U)^{-1}U^{\ast}S^{-1}$
$=S^{-1}-S^{-1}UA^{-1}U^{\ast}S^{-1}$
@@
$\widetilde{Y}_{n}=\begin{cases}\gamma^{n-1}U+Z_{n},&\text{for }n\leq p;\\[5.0pt] d_{n}U+W_{n},&\text{for }n\geq p+1,\end{cases}$ (55)
where
@@
Thus, the mutual information between $V_{1}^{n}$ and $Y_{1}^{n}$ can be expressed as
@@
$=-\frac{1}{2n}\log\mathbb{E}\left[\left|U-\mathbb{E}[U|\widetilde{Y}_{1}^{n}\right|^{2}\right].$
Define $G=\begin{bmatrix}1&\gamma&\cdots&\gamma^{p-1}\end{bmatrix}^{T}$, $D_{n}=\begin{bmatrix}d_{p+1}&d_{p+2}&\cdots&d_{n}\end{bmatrix}^{T}$ and $\widetilde{D}_{n}=\begin{bmatrix}G^{T}&D_{n}^{T}\end{bmatrix}^{T}$. Since $U$ and $\widetilde{Y}_{1}^{n}$ are jointly Gaussian, letting $\Sigma_{Z}=\mathrm{Cov}(Z_{1}^{p},(Z_{1}^{p})^{T})$, we obtain
$\mathbb{E}[|U-\mathbb{E}[U|\widetilde{Y}_{1}^{n}|^{2}]$ $=1-\begin{bmatrix}G^{T}&D_{n}^{T}\end{bmatrix}\begin{bmatrix}\Sigma_{Z}+GG^{T}&GD_{n}^{T}\\[2.0pt] D_{n}G^{T}&I+D_{n}D_{n}^{T}\end{bmatrix}^{-1}\begin{bmatrix}G\\[2.0pt] D_{n}\end{bmatrix}$ (58)
$=1-\widetilde{D}_{n}^{T}\left(\begin{bmatrix}\Sigma_{Z}&0\\[2.0pt] 0&I\end{bmatrix}+\widetilde{D}_{n}\widetilde{D}_{n}^{T}\right)^{-1}\widetilde{D}_{n}$
$\overset{(a)}{=}\frac{1}{1+G^{T}\Sigma_{Z}^{-1}G+D_{n}^{T}D_{n}}$
$=\frac{1}{1+G^{T}\Sigma_{Z}^{-1}G-L^{2}(\gamma^{-1})G^{T}G+\frac{L^{2}(\gamma^{-1})}{\gamma^{2}-1}(\gamma^{2n}-1)},$
where $(a)$ follows from Lemma 3.1. Substituting this result into (57) and taking the limit as $n\to\infty$, we obtain
$\lim_{n\to\infty}\frac{1}{n}I(V_{1}^{n};Y_{1}^{n})=\log|\gamma|.$ (59)
Moreover, following a similar approach as in (27), we have
@@
$=\mathbb{E}[|V_{n}-\mathbb{E}[V_{n}|\widetilde{Y}_{1}^{n-1}]|^{2}]$
$=\gamma^{2(n-1)}\mathbb{E}[|U-\mathbb{E}[U|\widetilde{Y}_{1}^{n-1}]|^{2}].$
Substituting (58) into (60) and letting $n\to\infty$ yields
$\lim_{n\to\infty}\mathbb{E}[X_{n}^{2}]$ $=\frac{\gamma^{2}-1}{L^{2}(\gamma^{-1})}$ (61)
@@
Consequently, to determine $C_{\mathrm{SK1}}(P)$, one must maximize the mutual information rate in (59) subject to the power constraint (61). This completes the proof, as desired.